OpenStudy (anonymous):

Just going to throw these up here i'm not really looking for answer really just the formulas if you want to give answer that's fine as well 18. A 1.6 m wave has a frequency of 355 Hz. What is its speed? 19. An object with a charge of -4.0 µC is 0.45 m from a second object and experiences an repulsive force of 1.80 N. What is the magnitude of the charge on the second object? 20. The large window air conditioner in Anita Breeze's room draws 11 amps of current. Determine the quantity of charge that will pass through Anita's window AC in the next 10 minutes.

OpenStudy (anonymous):

21. A coffee cup immersion heater utilizes a heating coil with a resistance of 8.5 Ω. Determine the current through the coil when operated at 110 V. 22. The power of a 1.5-volt alkaline cell varies with the number of hours of operation. A brand new D-cell can deliver as much as 13 A through a copper wire connected between terminals. Determine the power of a brand new D-cell. 23. A 5.0 m portion of wire carries a current of 8.0 A from east to west. It experiences a magnetic field of 6.0 × 10–4 T running from south to north. What is the magnitude of the magnetic force on the wire? 24. A proton (with charge of 1.6 x 10-19 C) moves perpendicularly at 5.0 × 105 m/s through a magnetic field that points upward and has a magnitude of 6.5 × 10–4 T. What is the magnitude of the magnetic force on the proton?

OpenStudy (anonymous):

18 : \[V = f \lambda\]

OpenStudy (anonymous):

\[\lambda = \frac{ c }{ f }\] frequence

OpenStudy (anonymous):

Q.19. \[F =k \frac{ q _{1}q _{2} }{ r ^{2} }\]

OpenStudy (anonymous):

Where k is the constant. q1 the charge of first object, q2 is charge of 2nd object. r is the distance between them

OpenStudy (anonymous):

Q.20. \[I = \frac{ q }{ t }\]

OpenStudy (anonymous):

and \[q=It\]

OpenStudy (anonymous):

21. U= R.I , i is the current

OpenStudy (anonymous):

Q.22. Power = VI

OpenStudy (anonymous):

Q.23. \[F = ILB \sin \Theta\]

OpenStudy (anonymous):

where " theta" is the angle between conductor and field.

OpenStudy (anonymous):

Q.24 correction \[F=qVB \sin \theta \]